HelpMePlz:

A wheel of radius 25 ft. is rotating 13 RPM counterclockwise. Considering a point on the rim of the rotating wheel, what is the angular speed ω in rad/sec and the linear speed v in ft/sec?

Champion:

Hints: 1 RPM=1 rotation / minute = 1*π radians / 60 sec. = π/60 rad/sec. Linear speed at the rim is the tangential speed, given by v=rω, where r=radius, and ω=angular velocity.